3. A 13.0 % solution of K2CO3 by mass has a density of 1.09 g/cm³ . Calculate the molarity of the solution.




1
Expert's answer
2021-11-29T04:37:33-0500

Let the solution be of mass 100 g

So the amount of K2CO3 will be 13% of 100 g i.e. 13 g

Molar mass of K2CO3 = 138.205 g /mol

n(K2CO3) "= \frac{13 }{ 138.205} = 0.094 \; moles"

"Density = \frac{mass }{ volume}"

Volume of the solution "= \frac{mass }{ density}"

"= \frac{100 }{1.09} \\

= 91.74 \; mL \\

Molarity = \frac{no. \;of \;moles }{ volume \;of \;solution\; in \;L} \\

M = \frac{0.094 }{ 0.09174} \\

M = 1.024"

The molarity of the solution is 1.024 M